OpenStudy (anonymous):

can some one explain how 25(4a^2-4ab+b^2) = 25(2a-b)(2a-b) when factoring completely?

OpenStudy (agreene):

to be specific, 25(2a-b)^2 is the lowest way to factor this. Because: 4a^2-4ab+b^2 factors to (2a-b)(2a-b) = (2a-b)^2 We then add in our constant multiplier and have 25(2a-b)^2

OpenStudy (anonymous):

i understand that 25(2a-b)(2a-b) = 25(2a-b)^2 i just dont understand how you get from 4a^2-4ab+b^2 to (2a-b)^2

OpenStudy (anonymous):

thats what i dont understand..

OpenStudy (agreene):

4a^2-4ab+b^2 2a(2a-b+b^2) (2a-b)(2a-b) does that help?

OpenStudy (anonymous):

uhhhmm i dont think so :/ im having a hard time trying to grasp this.

OpenStudy (anonymous):

omg college algebra! im in the 10th grade! haha thanks though!